# A Generalized Theory for Optical Cooling of a Trapped Atom with Spin

### v2

@@ IV.2 Spin cooling theory
Next we look at GM for $F=1$ to $F^{\prime}=1$. The CG coefficients are in Fig. 5(c). One optical selection rule is that the $|F,m=0\rangle$ to $|F^{\prime},m=0\rangle$ is forbidden, due to the photon needing to add one unit of angular momentum. Because of this dark state, all the population pumps to $m=0$. But in contrast to PG, this state is now dark. For that reason GM requires blue-detuned light ($\Delta>0$) to make this state the lowest energy. The simulation in Fig. 5(d) again shows that optimal cooling occurs when the light shift equals the trapping frequency. The population however drops close to ground-state. The optimal cooling is plotted in Fig. 5(e), along with the simple model with effective parameters $\gamma_{-}/\gamma_{+}=7.5\times 10^{-4}$ and heating scattering $\gamma_{h}/\gamma_{+}=1.5$. For this large optical pumping asymmetry where the reservoir state is dark, such as found in GM and EIT cooling, $\gamma_{\scriptscriptstyle-}=0$, and $s\approx\frac{\gamma_{\scriptscriptstyle+}}{\eta^{2}\gamma_{h}}$. In this dark cooling limit, the final population is instead limited by $\eta$.
@@ IV.3 $\Lambda$-GM and EIT Cooling
Figure 7: (a) EIT and $\Lambda$-enhanced GM decrease temperature by establishing more dark states to cool into. (b) EIT cooling uses two lasers with Rabi frequencies $\Omega_{1}$ and $\Omega_{2}$ to drive a two-photon transition from $|B,n\rangle$ to $|D,n-1\rangle$. (c) Simulations of population versus the two-photon detuning. Best cooling occurs at the two-photon resonance when $\Delta_{1}=\Delta_{2}$. A heating resonance occurs to the right. Parameters are $\eta=0.05$, $\Delta=10\Gamma$, and $\Omega_{1}=\Omega_{2}$. The cooling power is set so that the light shift equals the trapping frequency. The inset shows that as the cooling power is lowered, the resonance shifts over. When the light shift becomes small compared to the trapping frequency, the two-photon detuning must compensate to drive $n\rightarrow n-1$. (d) Comparison of GM and $\Lambda$-GM. Light dashed is simple model from Eq. 73 with parameters $\gamma_{h}/\gamma_{+}=2.0$ and $\gamma_{-}/\gamma_{+}=0.006$ for $\Lambda$-GM, and $\gamma_{s}/\gamma_{+}=5.0$ and $\gamma_{+}/\gamma_{-}=0.11$ for GM.
@@
The simple models reveal a crucial insight: lower populations are achieved when pumping to darker states. Gray molasses (GM) cooling generates dark states through angular momentum selection rules. Electromagnetically induced transparency (EIT) cooling and $\Lambda$-type gray molasses ($\Lambda$-GM) create additional dark states by incorporating multiple ground-states addressed with a second laser frequency in an EIT configuration, as illustrated in Fig. 7(a). These dark states arise from the interference of two lasers coupling the ground-states to a common excited state.
@@
Historically, the term “EIT cooling” has been primarily associated with trapped ions [63, 64], where the two ground-states are typically different Zeeman sublevels. In contrast, “$\Lambda$-GM” is commonly used in the context of neutral atoms [20, 61], where the two ground-states belong to different hyperfine manifolds of the ground-state.
@@
Cooling to the motional ground-state using EIT has been theoretically predicted and achieved experimentally in $Ca^{+}$ ion using Zeeman sublevels of the D1 line [56, 63]. $\Lambda$-enhanced GM cooling on the other hand is proven effective not only for bound atoms [21] but also for cooling and imaging molecules in optical traps [65].
@@
EIT addresses two ground-states $|g_{1}\rangle$ and $|g_{2}\rangle$ with two different Rabi frequencies $\Omega_{1}$ and $\Omega_{2}$. These two states are typically from different hyperfine ground-states. Each ground-state is addressed by a laser frequency to a single excited state with corresponding detunings $\Delta_{1}$ and $\Delta_{2}$. The Hamiltonian is written with $g_{1}$ in the rotating frame of laser one and $g_{2}$ in the rotating frame of laser two,
@@
We have two dipole operators, $\hat{D}_{1}=|g_{1}\rangle\langle e|$ and $\hat{D}_{2}=|g_{2}\rangle\langle e|$. The interaction Hamiltonian is
@@
The two collapse operators represent the spontaneous emission of the single excited state to both ground-states,
@@
Next, we perform a unitary transformation of the two ground-states into two superpositions so that only one of them is coupled to the two lasers, and the other is the EIT dark state, shown in Fig. 7(b). The dark state is dark due to the destructive interference of the population in the excited state from the two ground-states. The basis transformation is
@@
where $\Omega_{\text{rms}}=\sqrt{\Omega_{1}^{2}+\Omega_{2}^{2}}$.
Under this transformation, the bare Hamiltonian is
$H_{0}=\nu a^{\dagger}a+\Delta_{\scriptscriptstyle B}|g_{\scriptscriptstyle B}\rangle\langle g_{\scriptscriptstyle B}|+\Delta_{\scriptscriptstyle D}|g_{\scriptscriptstyle D}\rangle\langle g_{\scriptscriptstyle D}|$
$+\Omega_{\scriptscriptstyle BD}\bigl{(}|g_{\scriptscriptstyle B}\rangle\langle g_{\scriptscriptstyle D}|+|g_{\scriptscriptstyle D}\rangle\langle g_{\scriptscriptstyle B}|\bigr{)},$ (78)
with the energies of the bright and dark states as
@@
$\Delta_{\scriptscriptstyle D}=\left(\Omega_{1}^{2}\,\,\Delta_{2}+\Omega_{1}^{2}\,\,\Delta_{2}\right)/\Omega_{\text{rms}}^{2},$ (79)
and a coupling between the dark and bright states
@@
$H_{I}=\frac{1}{2\Omega_{\text{rms}}}\left[D_{B}^{\dagger}\,\left(\Omega_{1}^{2}\,e^{ik_{1}\hat{x}}+\Omega_{2}^{2}\,e^{ik_{2}\hat{x}}\right)\right.$
$\left.+D_{D}^{\dagger}\Omega_{1}\Omega_{2}\,\left(e^{ik_{1}\hat{x}}-e^{ik_{2}\hat{x}}\right)+h.c.\right].$ (81)
In the case where the two lasers are two-photon resonant, the Rabi frequencies are equal $\Omega_{1}=\Omega_{2}$, and the two lasers are counter-propagating $k_{1}=-k_{2}$, the expressions takes a simple form
$H_{0}=\nu a^{\dagger}a+\Delta\left(|g_{\scriptscriptstyle B}\rangle\langle g_{\scriptscriptstyle B}|+|g_{\scriptscriptstyle D}\rangle\langle g_{\scriptscriptstyle D}|\right)$ (82)
$H_{I}=\Omega\left(\cos(k\hat{x})D_{B}^{\dagger}+\sin(k\hat{x})D_{D}^{\dagger}\right)$ (83)
$\approx\Omega\left(D_{B}^{\dagger}+k\hat{x}\,\,D_{D}^{\dagger}\right).$ (84)
In the LD regime, only the bright state is coupled to the excited state. To the zeroth order, the dark state is not coupled. But the dark state can be driven directly only after changing the motional state. This Hamiltonian now looks similar to the spin cooling case.
An interesting case is if, instead, the lasers are co-propagating with $k_{1}=k_{2}$, then $H_{I}=\Omega D_{B}^{\dagger}e^{ikx}$, and the dark state is dark even to $n$ changing transitions. Co-propagating light cannot change the momentum, prohibiting cooling.
@@
Next, we can adiabatically eliminate the excited state using Eq. 15 and 16 to get the effective ground-state dynamics. In the original basis, the effective ground-state Hamiltonian is
@@
$H_{\text{eff}}=\nu\hat{a}^{\dagger}\hat{a}(\Delta_{1}-\frac{\Omega_{1}^{2}}{\Delta})|g_{1}\rangle\langle g_{1}|+(\Delta_{2}-\frac{\Omega_{2}^{2}}{\Delta})|g_{2}\rangle\langle g_{2}|$
@@
In the bright and dark basis with $\Omega_{1}=\Omega_{2}$ and $k_{2}=-k_{1}$, the effective ground-state Hamiltonian is
$H_{\text{eff}}=\nu a^{\dagger}a+\frac{\Omega_{1}^{2}+\Omega_{2}^{2}}{2\Delta}|g_{\scriptscriptstyle B}\rangle\langle g_{\scriptscriptstyle B}|$
$+\left(ik\hat{x}\frac{\Omega_{1}\Omega_{2}}{\Delta}|g_{\scriptscriptstyle B}\rangle\langle g_{\scriptscriptstyle D}|+h.c.\right)$ (86)
and the collapse operators are
$L_{\text{eff},\pm,i}=\frac{\sqrt{\Gamma}\,\Omega_{\text{rms}}}{\Delta-i\Gamma/2}\,\hat{D}_{i}\left(\,\hat{D}_{B}\pm ik\hat{x}(\,\hat{D}_{B}+\frac{2\Omega_{1}\Omega_{2}}{\Omega^{2}_{\text{rms}}}\,\hat{D}_{D})\right).$ (87)
We can now see that the Hamiltonian and collapse operator take a similar form as before. The position-independent Hamiltonian results in a light shift for only the bright state. Therefore, just like GM, EIT also works in the blue-detuned regime so that the dark state is lower energy. The position-dependent Hamiltonian results in coupling between neighboring motional states. The collapse operator also has a position-independent term, which optically pumps from the bright state to the dark state, and a position-dependent term from recoil heating. So if you are at two photon resonance condition ($\Lambda$-enhancement) in GM, the formalism and the effectiveness of $\Lambda$-GM is identical to EIT.
In Fig. 7(c), we plot the cooling performance as a function of detuning. In the main plot, the light shift is equal to the trapping frequency, so the optimal cooling occurs when $\Delta_{1}=\Delta_{2}$. The inset shows the cooling performance as the cooling power is decreased. The relative detuning can be adjusted to bring the system back into resonance and recover the cooling efficiency. Fig. 7(d) compares the optimal cooling performance for the $F=1/2,3/2$ to $F^{\prime}=3/2$ system, with and without the $\Lambda$-enhancement (i.e., with and without the $F=1/2$ state). The inclusion of the $F=1/2$ state in the $\Lambda$-configuration significantly improves the cooling performance. The simulation results are also fitted to the theoretical model to extract the effective cooling parameters.
IV.4 Raman-sideband cooling
Figure 8: Raman-sideband cooling. The first step is a Raman transition between spin states and decreasing the motional state by 1. The second step is spontaneous emission back to original spin, which keeps atom in $n-1$ in the Lamb-Dicke regime. The cycle repeats until the population is in the ground-state.
Raman sideband (RSB) cooling has achieved the highest ground-state populations in neutral atoms [66, 22, 67], ions [27], optical tweezers [24, 25, 26], and molecules [68, 69]. The working principle of RSB cooling is illustrated in Fig. 8(a).
In RSB cooling, two coherent, far-detuned lasers drive the transition from $|g_{1},n\rangle$ to $|g_{2},n-1\rangle$, where $|g_{1}\rangle$ and $|g_{2}\rangle$ are two ground-states, and $n$ represents the motional quantum number. The light shifts and scattering rates from these cooling lasers are typically negligible due to the large detuning. A third laser, tuned closer to resonance, performs optical pumping and exploits selection rules to ensure that the dark state remains dark.
The effective Raman coupling strength between the two ground-states is given by $\Omega_{1}\Omega_{2}/\Delta$, where $\Omega_{1}$ and $\Omega_{2}$ are the Rabi frequencies of the two cooling lasers, and $\Delta$ is the detuning from the excited state. The effective ground-state Hamiltonian for RSB cooling is:
@@ IV.4 Raman-sideband cooling
where $\hat{x}$ is the position operator, and $k_{1}$ and $k_{2}$ are the wave vectors of the two cooling lasers. By carefully tuning the laser frequencies and polarizations, this Raman coupling can be engineered to drive transitions that reduce the motional quantum number, leading to cooling.
@@
RSB cooling addresses this limitation by increasing the detuning of the cooling lasers until the collapse operator $L$ becomes negligible, effectively eliminating spontaneous emission and repumping from the cooling lasers. The additional repump beam, optimized using angular momentum selection rules, ensures that the state $|g_{1}\rangle$ remains dark to it, making the dark state immune to position-dependent spontaneous emission.
@@
While RSB cooling offers significant advantages, its fundamental limitations are often technical rather than theoretical. The coherence time between the two ground-states is typically limited to a few milliseconds due to factors such as magnetic field fluctuations or laser decoherence. For efficient cooling, the Raman Rabi frequency $\Omega_{R}$ should exceed 10 kHz. However, the Rabi frequency for the $n$ to $n$ transition is larger than that of the $n$ to $n-1$ transition by a factor of $1/\eta$, resulting in both transitions being driven simultaneously. Further analysis is needed to determine the cooling limit in the presence of these technical constraints.
While RSB cooling offers a pathway to overcome the limitations of EIT and $\Lambda$-GM cooling schemes, its implementation is more complex experimentally. However, it potentially enables more efficient ground-state cooling in a wider range of atomic and molecular systems.

### v3

@@ IV.2 Spin cooling theory
Next we look at GM for $F=1$ to $F^{\prime}=1$. The CG coefficients are in Fig. 6(c). One optical selection rule is that the $|F,m=0\rangle$ to $|F^{\prime},m=0\rangle$ is forbidden, due to the photon needing to add one unit of angular momentum. Because of this dark state, all the population pumps to $m=0$. But in contrast to PG, this state is now dark. For that reason GM requires blue-detuned light ($\Delta>0$) to make this state the lowest energy. The simulation in Fig. 6(d) again shows that optimal cooling occurs when the light shift equals the trapping frequency. The population however drops close to ground state. The optimal cooling is plotted in Fig. 6(e), along with the simple model with effective parameters $\gamma_{-}/\gamma_{+}=7.5\times 10^{-4}$ and heating scattering $\gamma_{h}/\gamma_{+}=1.5$. For this large optical pumping asymmetry where the reservoir state is dark, such as found in GM and EIT cooling, $\gamma_{\scriptscriptstyle-}=0$, and $s\approx\frac{\gamma_{\scriptscriptstyle+}}{\eta^{2}\gamma_{h}}$. In this dark cooling limit, the final population is instead limited by $\eta$.
@@ IV.3 $\Lambda$-GM and EIT Cooling
Figure 8: (a) EIT and $\Lambda$-enhanced GM decrease temperature by creating dark states from superpositions of ground states using two coherent lasers. (b) EIT cooling uses two lasers with Rabi frequencies $\Omega_{1}$ and $\Omega_{2}$ to address the two ground states, driving a two-photon transition from $|B,n\rangle$ to $|D,n-1\rangle$. (c) Master equation simulations of population versus the two-photon detuning. Best cooling occurs at the two-photon resonance when $\Delta_{1}=\Delta_{2}$. A heating resonance occurs to the right. Parameters are $\eta=0.05$, $\Delta=10\Gamma$, $\nu=0.05\,\Gamma$, $\Omega_{1}=\Omega_{2}$, and 8 harmonic states. The cooling power is set so that the light shift equals the trapping frequency. The inset shows that as the cooling power is lowered, the resonance shifts over. When the light shift becomes small compared to the trapping frequency, the two-photon detuning must compensate to drive $n\rightarrow n-1$. (d) Comparison of GM and $\Lambda$-GM. The $\Lambda$-GM model contains a $F=1/2$ and $F=3/2$ ground state with two laser frequencies. The GM model just has a single ground state $F=3/2$. Light dashed is simple model from Eq. 72 with effective parameters.
@@
The spin models reveal two strategies for lowering the temperature: first, to make the lower energy state dark to optical pumping back into the other spin state, and second to decrease the recoil heating from spontaneous emission. In terms of the collapse operator, these correspond to the position-independent and position-dependent terms. Gray molasses (GM) cooling generates dark states through optical angular momentum selection rules, for example the $m_{F}=0$ to $m_{F}^{\prime}=0$ state is dark for $F^{\prime}=F$. Electromagnetically induced transparency (EIT) cooling and $\Lambda$-type gray molasses ($\Lambda$-GM) create dark states by creating superpositions of multiple ground states in an EIT configuration that are dark to the scattering. The second set of ground states are addressed with a second coherent laser, as illustrated in Fig. 8(a). These dark states can be viewed as arising from the interference of the two lasers coupling the ground states to a common excited state.
@@
Historically, the term “EIT cooling” has been primarily associated with trapped ions [66, 67], where the two ground states are typically different Zeeman sublevels. In contrast, “$\Lambda$-GM” is commonly used in the context of neutral atoms [20, 62, 21], where the two ground states belong to different hyperfine manifolds of the ground state. $\Lambda$-GM modifies conventional GM with a second coherent laser to address the different hyperfine states, creating dark states with multiple spin states. EIT cooling often instead refers to optical pumping to a single spin state and only using two ground spin states.
@@
Cooling to the motional ground state using EIT has been theoretically predicted and achieved experimentally in Ca ion using Zeeman sublevels of the D1 line [56, 66]. $\Lambda$-enhanced GM cooling on the other hand is proven effective not only for bound atoms [21] but also for cooling and imaging molecules in optical traps [68].
To model $\Lambda$-GM or EIT cooling, we use the standard EIT formalism, for example in Ref. [69], to transform the system to the bright and dark basis. Then we apply the formalism from this paper to show that it produces similar position dependent and independent operators as in the spin cooling model. We will then discuss the fundamental limits of this type of cooling.
@@
EIT addresses two ground states $|g_{1}\rangle$ and $|g_{2}\rangle$ with two different Rabi frequencies $\Omega_{1}$ and $\Omega_{2}$, as shown in Fig. 8(a). These two states are typically from different hyperfine ground states. Each ground state is addressed by a laser frequency to a single excited state with corresponding detunings $\Delta_{1}$ and $\Delta_{2}$. The Hamiltonian is written with $g_{1}$ in the rotating frame of laser one and $g_{2}$ in the rotating frame of laser two,
@@
We have two dipole lowering operators, $\hat{D}_{1}=|g_{1}\rangle\langle e|$ and $\hat{D}_{2}=|g_{2}\rangle\langle e|$. The interaction Hamiltonian is
@@
The two collapse operators represent the spontaneous emission of the single excited state to both ground states,
@@
Next we transform the ground states into the bright and dark basis. We perform a unitary transformation of the two ground states into two superpositions so that only one of them is coupled to the two lasers, and the other is not. The superposition that is dark to the light is the EIT dark state, shown in Fig. 8(a). The dark state is dark due to the destructive interference of the population in the excited state from the two ground states.
The basis transformation to the bright and dark basis is
@@
where $\Omega_{\text{rms}}=\sqrt{\Omega_{1}^{2}+\Omega_{2}^{2}}$. Under this unitary transformation, the bare Hamiltonian becomes
$\tilde{H}_{0}=\nu a^{\dagger}a+\Delta_{\scriptscriptstyle B}|g_{\scriptscriptstyle B}\rangle\langle g_{\scriptscriptstyle B}|+\Delta_{\scriptscriptstyle D}|g_{\scriptscriptstyle D}\rangle\langle g_{\scriptscriptstyle D}|$
$+\Omega_{\scriptscriptstyle BD}\bigl{(}|g_{\scriptscriptstyle B}\rangle\langle g_{\scriptscriptstyle D}|+|g_{\scriptscriptstyle D}\rangle\langle g_{\scriptscriptstyle B}|\bigr{)}.$ (77)
The energies of the bright and dark states are
@@
$\Delta_{\scriptscriptstyle D}=\left(\Omega_{2}^{2}\,\,\Delta_{1}+\Omega_{1}^{2}\,\,\Delta_{2}\right)/\Omega_{\text{rms}}^{2},$ (78)
and the coupling between the dark and bright states is
@@
When $\Delta_{1}=\Delta_{2}$, both states are detuned by the same amount from the excited state and the Raman transition is resonant. Then the coupling in the bare Hamiltonian is $\Omega_{BD}=0$.
@@
$\tilde{H}_{I}=\frac{1}{2\Omega_{\text{rms}}}\left[\hat{D}_{B}^{\dagger}\,\left(\Omega_{1}^{2}\,e^{ik_{1}\hat{x}}+\Omega_{2}^{2}\,e^{ik_{2}\hat{x}}\right)\right.$
$\left.+\hat{D}_{D}^{\dagger}\Omega_{1}\Omega_{2}\,\left(e^{ik_{2}\hat{x}}-e^{ik_{1}\hat{x}}\right)+h.c.\right].$ (80)
If we assume that the two lasers are counter-propagating $-k_{1}=k_{2}=k$, assume the Rabi frequencies are the same, and expand the interaction Hamiltonian to first order in the LD parameter, we get
$\tilde{H}_{I}=\frac{\Omega}{\sqrt{2}}\left(\hat{D}_{B}^{\dagger}+ikx\hat{D}_{D}^{\dagger}+h.c.\right)$ (81)
From this form, we can draw some interesting conclusions. If the atom is not oscillating, only the bright state is coupled to the excited state. The dark is state is completely dark and population will quickly optically pump to it. When the atom oscillates, the dark state is driven to one of the sidebands.
@@
Next, we can adiabatically eliminate the excited state using Eq. 16 and 17 to get the effective ground state dynamics. In the bright and dark basis, the effective ground state Hamiltonian is
@@
$\tilde{H}_{\text{eff}}=\nu a^{\dagger}a+\frac{\Omega_{1}^{2}+\Omega_{2}^{2}}{2\Delta}|g_{\scriptscriptstyle B}\rangle\langle g_{\scriptscriptstyle B}|$
$+\left(ik\hat{x}\,\Omega_{r}\,|g_{\scriptscriptstyle B}\rangle\langle g_{\scriptscriptstyle D}|+h.c.\right)$ (82)
and the collapse operators are
$\tilde{L}_{\text{eff},\pm,i}=\sqrt{\gamma_{r}}\,\hat{D}_{i}\left(\,\hat{D}^{\dagger}_{B}\pm ik\hat{x}(\,\hat{D}_{B}^{\dagger}+\frac{2\Omega_{1}\Omega_{2}}{\Omega^{2}_{\text{rms}}}\,\hat{D}^{\dagger}_{D})\right),$ (83)
where $i=1,2$ are the two ground states, and the off-resonant scattering rate from both beams is
$\gamma_{r}=\Gamma\,\frac{\Omega_{1}^{2}+\Omega_{2}^{2}}{\Delta^{2}},$ (84)
and the Raman Rabi rate between the bright and dark state is
$\Omega_{r}=\frac{\Omega_{1}\Omega_{2}}{2\Delta}$ (85)
The form of the operators are similar to the GM spin cooling model. In the position independent effective Hamiltonian, the dark state is not driven and has no light shift. That makes it similar to the GM case, requiring blue-detuned light to make the bright state higher in energy. Similarly, the position independent collapse operators only couple the bright state, with an off-resonant scattering rate for $\Delta\gg\Gamma$. However the position dependent term couples the bright and dark state to different motional states, resulting in recoil heating at a rate $\approx\gamma_{s}\eta^{2}$. The effective Raman transition between $|g_{D},n\rangle$ and $|g_{B},n-1\rangle$ is $\Omega_{r}$. Just as we found in the spin cooling section, the transfer rate comes from the coherent coupling and the optical pumping, resulting in a transfer rate of $\Gamma_{r}\approx\Omega_{r}^{2}/\gamma_{r}$.
The picture for cooling is in Fig. 8(b). Population is pumped to the dark state. Then the two-photon process couples $|D,n\rangle$ to $|B,n-1\rangle$.
In Fig. 8(c), we simulate the cooling performance as a function of the two-photon detuning. These simulations are for the full ground and excited master equation without making the LD approximation. In the main plot, the bright state light shift is tuned to be equal to the trapping frequency, in which case optimal cooling occurs when $\Delta_{1}=\Delta_{2}$. The cooling has an EIT resonance at $\Delta_{1}=\Delta_{2}$. The linewidth of this resonance is given by the off-resonant scattering rate $\gamma_{s}$. There is a heating resonance nearby when $\Delta_{1}-\Delta_{2}=2\nu$.
The inset shows in Fig. 8(c) simulates the cooling performance as the cooling light power is decreased. When the light shift is not equal to the trapping frequency, the two-photon detuning can be adjusted to bring the system back into resonance and recover the cooling efficiency.
Fig. 8(d) compares the optimal cooling performance for the $F=1/2,3/2$ to $F^{\prime}=3/2$ system with and without the $\Lambda$-enhancement (i.e., with and without the $F=1/2$ state). The inclusion of the $F=1/2$ state in the $\Lambda$-configuration significantly improves the cooling performance. The simulation results are also fitted to the theoretical model to extract the effective cooling parameters. The light dashed lines are fitting the simple spin model from Eq. 72 with parameters $\gamma_{h}/\gamma_{+}=2.0$ and $\gamma_{-}/\gamma_{+}=0.006$ for $\Lambda$-GM, and $\gamma_{s}/\gamma_{+}=5.0$ and $\gamma_{-}/\gamma_{+}=0.11$ for GM.
One important requirement for EIT cooling is that the two lasers are coherent and at the same frequency. A non-zero two-photon detuning contributes to a bright-dark coupling in Eq. 77 with a rate $\Omega_{BM}\approx\Delta_{1}-\Delta_{2}$. As the phase between the two lasers slips, the bright and dark basis changes, also producing this coupling. The other bright-dark coupling in the effective Hamiltonian scales as $\eta\Omega_{r}$. For resonance $\Omega_{r}\approx\nu$, this means that $\Delta_{1}-\Delta_{2}$ must be less than $\eta\nu$ to be negligible. For typical trapping frequencies of 100 kHz, this implies that the lasers must be coherent to within $\sim 1$ kHz, which occurs naturally if the the two frequencies are generated from the same light using an EOM or AOM.
IV.4 Raman-sideband cooling
Figure 9: Raman-sideband cooling. The first step is a Raman transition between spin states and decreasing the motional state by 1. The second step is spontaneous emission back to original spin, which keeps atom in $n-1$ in the Lamb-Dicke regime. The cycle repeats until the population is in the ground state.
Raman sideband (RSB) cooling has high fidelity ground state preparation in neutral atoms [70, 22, 71], ions [27], optical tweezers [24, 25, 26], and molecules [72, 73]. The working principle of RSB cooling is illustrated in Fig. 9(a).
In RSB cooling, two coherent, far-detuned lasers drive the transition from $|g_{1},n\rangle$ to $|g_{2},n-1\rangle$, where $|g_{1}\rangle$ and $|g_{2}\rangle$ are two ground states, and $n$ represents the motional quantum number. The light shifts and scattering rates from these cooling lasers are typically negligible due to the large detuning. A third laser, tuned closer to resonance, performs optical pumping and exploits selection rules to ensure that the dark state remains dark.
For the RSB case, we instead derive an effective ground state basis in the original $g_{1}$, $g_{2}$ basis instead of the bright and dark basis. In the original basis, the effective ground state Hamiltonian is
$H_{\text{eff}}=\nu\hat{a}^{\dagger}\hat{a}+(\Delta_{1}-\frac{\Omega_{1}^{2}}{\Delta})|g_{1}\rangle\langle g_{1}|+(\Delta_{2}-\frac{\Omega_{2}^{2}}{\Delta})|g_{2}\rangle\langle g_{2}|$
@@ IV.4 Raman-sideband cooling
The light shifts $\Omega_{1}^{2}/\Delta$ and $\Omega_{2}^{2}/\Delta$ shift the two states. The effective Raman coupling strength between the two ground states is given by $\Omega_{R}=\Omega_{1}\Omega_{2}/\Delta$, where $\Omega_{1}$ and $\Omega_{2}$ are the Rabi frequencies of the two cooling lasers, and $\Delta$ is the detuning from the excited state.
The transition matrix elements between the two ground states are
@@
where $k_{1}$ and $k_{2}$ are the wave vectors of the two cooling lasers. By carefully tuning the laser frequencies and polarizations, this Raman coupling can be engineered to drive transitions that reduce the motional quantum number, leading to cooling.
@@
RSB cooling addresses this limitation by increasing the detuning of the cooling lasers until the collapse operator $L$ becomes negligible, effectively eliminating spontaneous emission and repumping from the cooling lasers. The additional repump beam, optimized using angular momentum selection rules, ensures that the state $|g_{1}\rangle$ remains dark to it, making the dark state immune to position-dependent spontaneous emission. PG, GM, and $\Lambda$ had a position dependent collapse operator that drove heating in the dark state. This is because they use counter-propagating light that coherently couple spins to also optically pump. RSB uses a separate optical pumping beam that is completely dark to the dark state.
@@
While RSB cooling offers significant advantages, its fundamental limitations are often technical rather than theoretical. The coherence time between the two ground states is typically limited to a few milliseconds due to factors such as magnetic field fluctuations or laser decoherence. For efficient cooling, the Raman Rabi frequency $\Omega_{R}$ should exceed 10 kHz. However, the Rabi frequency for the $n$ to $n$ transition is larger than that of the $n$ to $n-1$ transition by a factor of $1/\eta$, resulting in both transitions being driven simultaneously. Further analysis is needed to determine the cooling limit in the presence of these technical constraints.
While RSB cooling offers a pathway to overcome the limitations of EIT and $\Lambda$-GM cooling schemes, its implementation is more complex experimentally. However, it potentially enables more efficient ground state cooling in a wider range of atomic and molecular systems.
Similar to $\Lambda$-GM, RSB also requires phase coherence of the laser. It also requires that $\Omega_{r}$ is less than the trapping frequency. However, RSB is usually performed with a separate set of lasers than the MOT beams.